The acceleration of a particle is bt(b is constant). The particle start s with velocity u. The distance travelled by particle in time t is

Grade:11

1 Answers

Khimraj
6 years ago
Here acceleration is not constant so you can not apply equation of motion. So apply basic concept.
acceleration a = dv/dt = bt
dv = btdt
\int_{u}^{v}dv = b\int_{0}^{t}tdt
v -u = bt2/2
v = ds/dt = u + bt2/2
So ds = ( u + bt2/2)dt
integerating both side
s = ut + bt3/6.
Hope it clears.
If you like answer then please approve it.